OpenStudy (anonymous):

60% of a number is 90. What is 50% of that same number? A. 30 B. 45 C. 75 D. 150

OpenStudy (kc_kennylau):

Please edit the question and use proper spacing next time :D So we know that 60% of some number is 90. Let's set x to be that number :) So we have 60%x=90 And then, we see that 60% is equivalent to 0.6. Then, the equation transforms to 0.6x=90 Solve that equation by dividing both sides by 0.6, and you get x=90/0.6 Simplify that and you get x=150 50% of the same number can be expressed as 50%x. And remember that 50%=0.5 Therefore, 50%x=0.5x And substitute x=150 inside: 50%x=0.5x=0.5*150=75 Problem solved :D

OpenStudy (anonymous):

thx